I do not know how to solve these equations -9x+6=-x+4, and-9x+6= -1x+4 can you help me with these equations?

Answer by AlexxsDad(4)

Your two equations are the same: -x+4 is the exact same thing as -1x+4. The "1" is usually omitted; x is the same as 1x (because there's only one), and -x is the same as -1x. So we have:
-9x+6=-x+4
Strategy: put x's on one side and number on the other
6-4 = 9x-x (study that move)
2 = 8x
or if you prefer:
8x = 2 (they're equal; it doesn't matter what side of the equal sign they're on)
divide both sides by 8 so your x is identified:
x = 2/8, better known as x = 1/4